How much water should be added to 10g of acetic acid to give a hydrogen ion concentration of 10^-3 M(given:pKa=4.74)?

Grade:11

1 Answers

vikas askiitian expert
13 years ago

let mass of water added is V litre...

molarity = 10/60V =1/6V                            (molar mass of acetic acid is 60)

[H+] = (KaM)1/2

M = [H+]2 /Ka .............1

H+ = 10-3  &        pka = -logKa = 4.74

                         Ka = 1.8*10-5

now M = 10-6 / 1.8*10-5 =1/6V

      V = 3 L